# Tensor networks enable the calculation of turbulence probability distributions

### v1

@@ Discussion
Ensuring the MPS algorithm maintains both accuracy and efficiency requires carefully selecting $\chi$ (see Suppl. Figs. S1, S3). Figure 2 indicates that varying the Damköhler number does not significantly affect the accuracy, while increasing the SGS mixing-rate requires $\chi$ to in turn increase as $\chi\sim\text{poly}(C_{\Omega})$ for accuracy to be maintained. Setting $\chi$ excessively high is expensive due to the $O(\chi^{4}d\log M)$ asymptotic cost of the algorithm. However, for slower mixing-rates, high accuracy is achievable at very low $\chi$ even when the chemical reaction rates are high. For instance, at $C_{\Omega}=0.25$, $\text{Da}=1.5$, the algorithm is accurate with just $\chi=32$. This is equivalent to respective $\mathcal{O}(10^{6})$ and $\mathcal{O}(10^{3})$ factor reductions in memory and computational costs [Methods, Eq. (12) and Suppl. Sec. 2] compared to conventional FD schemes; allowing the time-evolution to be executed on a single CPU core in only a couple of hours, instead of days on a supercomputer.
@@
The results shown here are only an early indication of what is possible: there exists great scope for improvement in both the algorithm and its implementation. For example, employing the tensor-cross algorithm to perform element-wise multiplications might reduce the complexity of our scheme to $\sim\chi^{3}$ without significantly sacrificing accuracy. Furthermore, better optimised software running on specialised computing architectures will allow for much larger bond-dimensions and system sizes: we are currently simulating a $M^{d}=128^{5}$ $=2^{35}$ grid at $\chi=128$, while the current record is a quantum physics simulation on a grid-equivalent of $M^{d}=2^{400}$ at $\chi=32768$, performed on a tensor processing unit pod [43].
@@
We decided to use an MPS ansatz because it closely matches the structure of the PDF for this particular flow; other flows may have different PDFs for which alternative ansatze could be better suited [44]. Fortunately, there exists a rich and growing selection of TNs to pick from, each carrying their advantages and disadvantages. These range from 2D generalisations of MPSs [45], hierarchical networks [46, 47] and even networks that might some day leverage quantum hardware [48]. There is also the exciting prospect of catering the TN ansatz to the structure of the PDF in an automated manner [49]. Typically, more complex ansatze are able to encode solutions with higher accuracy at lower $\chi$, but are costlier to manipulate. Balancing such considerations while exploring alternative TN geometries for probabilistic turbulence simulations is a promising avenue of future investigation.
@@
Turbulence is just one example of a complex system; there are many others, ranging from biological organisms to financial markets [50]. These kind of systems exhibit chaotic and unpredictable dynamics that ultimately require statistical descriptions [51]. The most fundamental way of doing so is by modelling their PDFs. Yet, such PDFs are typically prohibitively high-dimensional (as displayed here for the case of turbulence) which has made solving their governing Fokker-Planck equations infeasible, until now. This work is a first demonstration in how the problem can be overcome via a simple TN. More advanced TN ansatze and algorithms will be developed in time, holding the promise of enabling large-scale probabilistic simulations both within the field of fluid dynamics, and beyond.
Methods
@@
$\begin{split}\braket{U_{1}}/u_{0}&=\cos{kx_{1}}\sin{kx_{2}}\sin{kx_{3}}\\ &-e^{-\frac{(x_{2}/l_{0}-1/2)^{2}+(x_{3}/l_{0}-1/2)^{2}}{2(1/6)^{2}}},\\ \braket{U_{2}}/u_{0}&=\sin{kx_{1}}\cos{kx_{2}}\sin{kx_{3}},\\ \braket{U_{3}}/u_{0}&=-2\sin{kx_{1}}\sin{kx_{2}}\cos{kx_{3}}.\end{split}$ (7)
@@
$f(t=0)=\frac{1}{\sqrt{2\pi}(1/8)}\left\{\begin{array}[]{ll}e^{-\frac{(\varphi_{1}-3/4)^{2}+(\varphi_{2}-1/4)^{2}}{2(1/8)^{2}}},&\frac{1}{4}\leq x_{1}<\frac{3}{4},\\ e^{-\frac{(\varphi_{1}-1/4)^{2}+(\varphi_{2}-3/4)^{2}}{2(1/8)^{2}}},&\mbox{otherwise,}\end{array}\right.$ (8)
@@
In non-reactive flows ($\text{Da}=0$), $S=0$ and Eq. (9) can be cheaply and accurately solved using a standard FD scheme to obtain a “numerically exact” $\braket{\Phi_{\alpha}}$ solution (in the sense that there is no truncation error in $\chi$, as explained in Methods, Error measures). This is used to check the accuracy of the MPS algorithm in Figs. 2a and 2b. It is not possible to obtain a numerically exact $\braket{\Phi_{\alpha}}$ when $\text{Da}>0$, because a closure model would be required for $\braket{S_{\alpha}}$.
Finite difference discretisation
The simulations are performed on equidistant Cartesian grids with $M=128$ gridpoints along each dimension. The derivatives in Eqs. (3) and (9) are discretised in a simple manner: the temporal derivative with an explicit Runge-Kutta 2 scheme, and a second-order-accurate central finite differences (CFD2) discretisation of the $\mathbf{x},\varphi_{1},\varphi_{2}$ derivatives.
@@
$\begin{split}&\frac{\Delta f}{\Delta t}+\braket{U_{i}}\frac{\Delta f}{\Delta x_{i}}-\frac{\Delta}{\Delta x_{i}}\left[(\gamma+\gamma_{\text{SGS}})\frac{\Delta f}{\Delta x_{i}}\right]=\\ &\frac{\Delta}{\Delta\varphi_{\alpha}}\left[\Omega_{\text{mix}}\left(\varphi_{\alpha}-\braket{\Phi_{\alpha}}\right)f+C_{\Omega}\mu\frac{\Delta f}{\Delta\varphi_{\alpha}}\right]\\ &-\frac{\Delta}{\Delta\varphi_{\alpha}}(S_{\alpha}f),\end{split}$ (10)
@@
Our MPS algorithm implements the aforementioned RK2-CFD2 scheme on the MPS manifold [14, 52]. This entails parameterising all the vectors (like $f$, $\braket{U_{i}}$ and $\braket{\Phi\alpha}$) in Eq. (10) as MPSs [38, Section 3.4], and the matrices (e.g. $\Delta/\Delta x_{i}$) as analogous matrix product operators (MPOs, [38, Section 3.5]). Then, within the MPS format, the time-stepping is performed in a standard manner using the arithmetic operations outlined in the Computational complexity section.
It is essential to control the bond-dimension during the MPS simulation. The arithmetic operations that time-evolve $f$ lead to its bond-dimension growing exponentially in time, if not truncated [38, Section 3.5.3]. In our code, we employ the singular values decomposition to truncate the bond-dimension of $f$ such that it is always limited to $\chi$. As for the other vectors and matrices, these objects remain constant in time and their bond-dimensions are all of order $\mathcal{O}(10)$.
@@
with $N=\log_{2}M^{d}$, ($M$ must be a power of $2$) being the number of tensors in the MPS, and $p(n)=\min(2^{n},2^{N-n},\chi)$ being the size of the $n$th bond of the MPS. The first sum gives the total number of parameters in the MPS, while the second sum represents the intrinsic gauge degrees of freedom of the MPS format [52]. When $\chi$ is maximal, i.e. $\chi=2^{\lfloor N/2\rfloor}$, we get $\text{NVPP}=2^{N}=M^{d}$ and that $f$ is represented exactly on the $M^{\times d}$ grid.
@@
with $\braket{\Phi_{\alpha}}(\chi)$ being extracted from the MPS-PDF solution of Eq. (10) while $\braket{\Phi_{\alpha}}(\text{exact})$ is the solution found by directly solving Eq. (9) with a standard RK2-CFD2 FD scheme; this solution is “numerically-exact” in the sense that it does not suffer from any truncation error in $\chi$ [although a truncation error from the FD discretisation itself remains, this error is slight due to the smoothness of $\braket{\Phi_{\alpha}}(\text{exact})$]. The RMSE functions are
@@
Here $\mathcal{E}_{t}$ is the temporally-averaged RMSE, and $\mathcal{E}_{t,\mathbf{x}}$ is averaged in both space and time.
@@
Acknowledgements
@@
NG acknowledges support by US-AFSOR grant FA8655-22-1-7027 and the EPSRC “Quantum Computing and Simulation Hub” (EP/P009565/1). PG acknowledges support by US-AFOSR grant FA9550-23-1-0014. DJ acknowledges support by the European Union’s Horizon Programme (HORIZON-CL42021-DIGITALEMERGING-02-10) Grant Agreement 101080085 QCFD, the Cluster of Excellence “Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG)- EXC 2056- project ID 390715994, and the Hamburg Quantum Computing Initiative (HQIC) project EFRE. The project is co-financed by ERDF of the European Union and by “Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB)”. The funders had no role in study design, data collection and analysis, decision to publish or preparation of the manuscript. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work (http://dx.doi.org/10.5281/zenodo.22558). Finally, we thank Professor Ard Louis at the University of Oxford for fruitful discussions.

### v2

@@ Discussion
Ensuring the MPS algorithm maintains both accuracy and efficiency requires carefully selecting $\chi$ (see Figs. S1 to S3). Figure 2 indicates that varying the Damköhler number does not significantly affect the accuracy, while increasing the SGS mixing-rate requires $\chi$ to in turn increase as $\chi\sim\text{poly}(C_{\Omega})$ for accuracy to be maintained. Setting $\chi$ excessively high is expensive due to the $O(\chi^{4}d\log M)$ asymptotic cost of the algorithm. However, for slower mixing-rates, high accuracy is achievable at very low $\chi$ even when the chemical reaction rates are high. For instance, at $C_{\Omega}=0.25$, $\text{Da}=1.5$, the algorithm is accurate with just $\chi=32$. This is equivalent to respective $\mathcal{O}(10^{6})$ and $\mathcal{O}(10^{3})$ factor reductions in memory and computational costs [Materials and Methods, Eq. (12) and Supplementary Text, Empirical computational cost] compared to conventional FD schemes; allowing the time-evolution to be executed on a single CPU core in only a couple of hours, instead of days on a supercomputer.
@@
The results shown here are only an early indication of what is possible: there exists great scope for improvement in both the algorithm and its implementation. For example, employing tensor-cross or other algorithms  (?) to perform element-wise multiplications might reduce the complexity of our scheme to $\sim\chi^{3}$ without significantly sacrificing accuracy. Furthermore, better optimised software running on specialised computing architectures will allow for much larger bond-dimensions and system sizes: we are currently simulating a $M^{d}=128^{5}$ $=2^{35}$ grid at $\chi=128$, while the current record is a quantum physics simulation on a grid-equivalent of $M^{d}=2^{400}$ at $\chi=32768$, performed on a tensor processing unit pod  (?).
@@
We decided to use an MPS ansatz because it closely matches the structure of the PDF for this particular flow; other flows may have different PDFs for which alternative ansatze could be better suited  (?). Fortunately, there exists a rich and growing selection of TNs to pick from, each carrying their advantages and disadvantages. These range from 2D generalisations of MPSs  (?), hierarchical networks  (?, ?) and even networks that might some day leverage quantum hardware  (?). There is also the exciting prospect of catering the TN ansatz to the structure of the PDF in an automated manner  (?). Typically, more complex ansatze are able to encode solutions with higher accuracy at lower $\chi$, but are costlier to manipulate. Balancing such considerations while exploring alternative TN geometries for probabilistic turbulence simulations is a promising avenue of future investigation.
@@
Turbulence is just one example of a complex system; there are many others, ranging from biological organisms to financial markets  (?). These kind of systems exhibit chaotic and unpredictable dynamics that ultimately require statistical descriptions  (?). The most fundamental way of doing so is by modelling their PDFs. Yet, such PDFs are typically prohibitively high-dimensional (as displayed here for the case of turbulence) which has made solving their governing Fokker-Planck equations infeasible, until now. This work is a first demonstration in how the problem can be overcome via a simple TN. More advanced TN ansatze and algorithms will be developed in time, holding the promise of enabling large-scale probabilistic simulations both within the field of fluid dynamics, and beyond.
Materials and Methods
@@
$\begin{split}\braket{U_{1}}/u_{0}&=\cos{kx_{1}}\sin{kx_{2}}\sin{kx_{3}}-e^{-\frac{(x_{2}/l_{0}-1/2)^{2}+(x_{3}/l_{0}-1/2)^{2}}{2(1/6)^{2}}},\\ \braket{U_{2}}/u_{0}&=\sin{kx_{1}}\cos{kx_{2}}\sin{kx_{3}},\\ \braket{U_{3}}/u_{0}&=-2\sin{kx_{1}}\sin{kx_{2}}\cos{kx_{3}}.\end{split}$ (7)
@@
$f(t=0)=\frac{1}{{2\pi}(1/8)^{2}}\left\{\begin{array}[]{ll}e^{-\frac{(\varphi_{1}-3/4)^{2}+(\varphi_{2}-1/4)^{2}}{2(1/8)^{2}}},&\frac{1}{4}\leq x_{1}<\frac{3}{4},\\ e^{-\frac{(\varphi_{1}-1/4)^{2}+(\varphi_{2}-3/4)^{2}}{2(1/8)^{2}}},&\mbox{otherwise,}\end{array}\right.$ (8)
@@
In non-reactive flows ($\text{Da}=0$), $S=0$ and Eq. (9) can be cheaply and accurately solved using a standard FD scheme to obtain a “numerically exact” $\braket{\Phi_{\alpha}}$ solution (in the sense that there is no truncation error in $\chi$, as explained in Materials and Methods, Error measures). This is used to check the accuracy of the MPS algorithm in Figs. 2a and 2b. It is not possible to obtain a numerically exact $\braket{\Phi_{\alpha}}$ when $\text{Da}>0$, because a closure model would be required for $\braket{S_{\alpha}}$.
Finite-difference discretisation
The simulations are performed on equidistant Cartesian grids with $M=128$ gridpoints along each dimension. The derivatives in Eqs. (3) and (9) are discretised in a simple manner: the temporal derivative with an explicit Runge-Kutta 2 scheme, and a second-order-accurate central finite-differences (CFD2) discretisation of the $\mathbf{x},\varphi_{1},\varphi_{2}$ derivatives.
@@
$\begin{split}&\frac{\Delta f}{\Delta t}+\braket{U_{i}}\frac{\Delta f}{\Delta x_{i}}-\frac{\Delta}{\Delta x_{i}}\left[(\gamma+\gamma_{\text{SGS}})\frac{\Delta f}{\Delta x_{i}}\right]=\\ &\frac{\Delta}{\Delta\varphi_{\alpha}}\left[\Omega_{\text{mix}}\left(\varphi_{\alpha}-\braket{\Phi_{\alpha}}\right)f+C_{\Omega}\mu\frac{\Delta f}{\Delta\varphi_{\alpha}}\right]-\frac{\Delta}{\Delta\varphi_{\alpha}}(S_{\alpha}f),\end{split}$ (10)
@@
Our MPS algorithm implements the aforementioned RK2-CFD2 scheme on the MPS manifold  (?, ?). This entails parameterising all the vectors (like $f$, $\braket{U_{i}}$ and $\braket{\Phi\alpha}$) in Eq. (10) as MPSs  (?, Section 3.4), and the matrices (e.g. $\Delta/\Delta x_{i}$) as analogous matrix product operators  (?, Section 3.5). Then, within the MPS format, the time-stepping is performed in a standard manner using the arithmetic operations outlined in the Computational complexity section.
It is essential to control the bond-dimension during the MPS simulation. The arithmetic operations that time-evolve $f$ lead to its bond-dimension growing exponentially in time, if not truncated  (?, Section 3.5.3). In our code, we employ the singular values decomposition to truncate the bond-dimension of $f$ such that it is always limited to $\chi$. As for the other vectors and matrices, these objects remain constant in time and their bond-dimensions are all of order $\mathcal{O}(10)$.
@@
with $N=\log_{2}M^{d}$, ($M$ must be a power of $2$) being the number of tensors in the MPS, and $p(n)=\min(2^{n},2^{N-n},\chi)$ being the size of the $n$th bond of the MPS. The first sum gives the total number of parameters in the MPS, while the second sum represents the intrinsic gauge degrees of freedom of the MPS format  (?). When $\chi$ is maximal, i.e. $\chi=2^{\lfloor N/2\rfloor}$, we get $\text{NVPP}=2^{N}=M^{d}$ and that $f$ is represented exactly on the $M^{\times d}$ grid.
@@
with $\braket{\Phi_{\alpha}}(\chi)$ being extracted from the MPS-PDF solution of Eq. (10) while $\braket{\Phi_{\alpha}}(\text{exact})$ is the solution found by directly solving Eq. (9) with a standard RK2-CFD2 FD scheme; this solution is “numerically-exact” in the sense that it does not suffer from any truncation error in $\chi$ [although a truncation error from the FD discretisation itself remains, this error is slight due to the smoothness of $\braket{\Phi_{\alpha}}(\text{exact})$]. The $\mathcal{E}_{t}$ function implements temporal-averaging, while $\mathcal{E}_{t,\mathbf{x}}$ performs averaging across space-time:
@@
Acknowledgments
We thank Professor Ard Louis at the University of Oxford for fruitful discussions and would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work (http://dx.doi.org/10.5281/zenodo.22558).
Funding:
@@ Acknowledgments
NG acknowledges support by US-AFSOR grant FA8655-22-1-7027 and the UKRI “Quantum Computing and Simulation Hub” grant EP/P009565/1. PG acknowledges support by US-AFOSR grant FA9550-23-1-0014. DJ acknowledges support by the European Union’s Horizon Programme (HORIZON-CL42021-DIGITALEMERGING-02-10) Grant Agreement 101080085 QCFD, the Cluster of Excellence “Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG)- EXC 2056- project ID 390715994, and the Hamburg Quantum Computing Initiative (HQIC) project EFRE. The project is co-financed by ERDF of the European Union and by “Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB)”.
Author contributions:
N.G. and P.G. conceived and planned the research project. N.G. formulated the matrix product state algorithm, did the analytical calculations and wrote the software. The numerical experiments were jointly designed by all the authors and executed by N.G. The numerical results were analysed and interpreted jointly by all the authors. N.G. and P. G. wrote the manuscript, with contributions from the other authors. The project was supervised by P. G. and S. B. Pope.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All software and data accompanying this manuscript are publicly available at https://github.com/nikitn2/tendeq and https://doi.org/10.5281/zenodo.14223424.
Supplementary Materials
Supplementary Text
Figs. S1 to S3
Reference (0)
Supplementary Materials for
Tensor networks enable the calculation of turbulence probability distributions
Nikita Gourianov, Peyman Givi, Dieter Jaksch, and Stephen B. Pope
Corresponding author. E-mail: nikgourianov@icloud.com
This PDF file includes:
Supplementary Text
Figures S1 to S3
Supplementary Text
Separability of Fokker-Planck equation
Here we explain how the Main Text Fokker Planck Equation (2) becomes separable when $C_{\Omega}=0$. Assume the PDF can be written in the form
$f(\varphi_{1},\varphi_{2},x_{1},x_{2},x_{3},t)=g(x_{1},t)h(\varphi_{1},t)R(\varphi_{2},x_{2},x_{3},t).$ (S1)
Inserting this into Equation (2), gives
$\begin{split}&hR\left(\frac{\partial g}{\partial t}+\braket{U_{1}}\frac{\partial g}{\partial x_{1}}-\frac{\partial}{\partial x_{1}}\left[\left(\gamma+\gamma_{\text{SGS}}\right)\frac{\partial g}{\partial x_{1}}\right]\right)+\\ &gR\left(\frac{\partial h}{\partial t}-\Omega_{\text{mix}}\frac{\partial}{\partial\varphi_{1}}\left[(\varphi_{1}-\braket{\Phi_{1}})h\right]-C_{r}\varphi_{2}\frac{\partial}{\partial\varphi_{1}}[\varphi_{1}h]\right)+\\ &gh\bigg{(}\frac{\partial R}{\partial t}+\sum_{i=2,3}\bigg{\{}\braket{U_{i}}\frac{\partial R}{\partial x_{i}}-\frac{\partial}{\partial x_{i}}\left[\left(\gamma+\gamma_{\text{SGS}}\right)\frac{\partial R}{\partial x_{i}}\right]\bigg{\}}-\\ &\Omega_{\text{mix}}\frac{\partial}{\partial\varphi_{2}}\left[(\varphi_{2}-\braket{\Phi_{2}})R\right]-C_{r}\varphi_{1}\frac{\partial}{\partial\varphi_{2}}[\varphi_{2}R]\bigg{)}=0.\end{split}$ (S2)
Note here that it would be possible to solve for $g,h,R$ separately if not for the nonlinear term $\braket{\Phi_{\alpha}}=\int\varphi_{\alpha}f\text{d}\varphi_{1}\text{d}\varphi_{2}$. This term can however be eliminated from the above equation by setting $C_{\Omega}=0\rightarrow\Omega_{\text{mix}}=0$, thus completely decoupling the solutions along $x_{1},\varphi_{1}$ from each other and the other dimensions. In fact, it is straightforward to show that setting $C_{\Omega}=0$ will decouple all the dimensions from each other, turning Eq. (2) of the Main Text into a separable PDE with product-state solutions. Thus at low $C_{\Omega}$, there should only be limited correlation between the different dimensions.
Empirical computational cost
The Main Text outlines the theoretical complexity of our algorithm. This section shows the empirical computational cost. The CPU-times referenced here are all measured on a single CPU core, by selecting the wall-time of the longest of the first $16$ time-steps (out of the full $M_{t}$ required to complete the time-evolution) of the algorithm (at $C_{\Omega}=1$, $\text{Da}=0$).
Single-core times as functions of $\chi,M$ are shown in Figs. S1 and S2. The first figure shows that at small $\chi$, the cost of the program is dominated by the overhead of the Python interpreter, while at intermediate $\chi$, the $\sim\chi^{4}$ element-wise products are too few to practically dominate the many $\sim\chi^{2}$ and $\sim\chi^{3}$ operations (outlined in the Main Text) of the algorithm. Fig. S2 empirically confirms the $\sim\log M$ scaling of the MPS algorithm, in contrast to the $\sim M^{d}$ cost required by an equivalent, standard FD scheme. Note how the MPS algorithm at $\chi=32,M=128$ requires $\mathcal{O}(1/10^{3})$ of the flops of the equivalent FD scheme.
Figure S1: Empirical cost of MPS-TN algorithm. The single-core CPU-time (in seconds) required for a typical time-step of the MPS algorithm is shown for different $\chi$. The algorithm was implemented in Python 3 [on top of the quimb library  (?)] and run on a single core from a Intel Xeon 8268 CPU.
Figure S2: MPS-TN versus FD scheme empirical cost. Equivalent single-core CPU-time to Fig. S1, except now plotted across $M$ for both the MPS algorithm at various $\chi$ and an equivalent, standard FD scheme (implemented in Python 3). For the FD scheme, the $M=128$ datapoint could not be measured due to the extreme memory load required.
Main Text Figs. 2b, 2c can be combined with Fig. S1 to show how the observed accuracy of the MPS algorithm increases with computational effort. This relationship is illustrated in Supplementary Fig. S3.
Figure S3: Accuracy convergence of MPS-TN algorithm versus empirical computational cost. Here the errors of Main Text Fig. 2 are plotted as functions of the single-core CPU time measured in Fig. S1. A corresponds to the Fig. 2b errors, and B to the Fig. 2c ones.